Use a standard algorithm to calculate :.7-2/7-1/6
1
Expert's answer
2019-09-03T10:23:54-0400

"7-\frac{2}{7}-\frac{1}{6}" "=\frac{7}{1}-\frac{2}{7}-\frac{1}{6}" "=\frac{(42\times7)-(6\times2)-(7\times 1)}{42}=\frac{275}{42}"